# Galactic Archaeology with Gaia

### v1

@@ 4.3 Pre-disc Milky Way
Luckily, additional chemical tags can be used to attempt to separate the accreted and in-situ born stars at [Fe/H]$<-1$. These ideas are first presented in Hawkins et al. (2015) who suggest that the accreted and in-situ halo stars attain distinct levels of [Al/Fe] abundance ratio. The stellar nucleosynthetic yield of aluminium (and sodium) is low at low metallicity but increases dramatically as more of C, N, O is pumped into the interstellar medium (see Kobayashi et al., 2006) which causes a delay in Al enrichment compared to conventional $\alpha$-elements. Compared to more massive systems, such as the progenitor of the Milky Way, dwarf galaxies are less efficient at retaining and re-using gas during star formation and are therefore slower to reach levels of metallicity when Al production becomes efficient. As a result, in a dwarf, by the time Al starts to be produced in earnest, type Ia SNe begin to dominate Fe production, thus inhibiting the growth of the [Al/Fe] ratio. In a more rapidly evolving Milky Way progenitor, Al production kicks in before the Type Ia SNe contribution rises and thus the young Galaxy can enjoy a period of over-abundance of Al relative to Fe. The systematic difference between [Al/Fe] levels in dwarf galaxies and the Milky Way proper is now well established (see e.g. Hasselquist et al., 2021) and is routinely taken advantage of to identify accreted stars in the nearby halo (e.g. Hawkins et al., 2015; Das et al., 2020; Horta et al., 2021).
@@ 4.3.1 Aurora
Instead of focusing on the accreted debris, Belokurov and Kravtsov (2022) combine [Al/Fe] measurements published as part of the APOGEE DR17 and the Gaia EDR3 astrometry to create a pure Milky Way in-situ sample, including stars at low metallicities (see top panel of Fig. 18). They show that while there may be some amount of false negatives, i.e. genuine in-situ stars identified as accreted, the in-situ selection is largely uncontaminated. Having access to a pure set of stars born only in the Milky Way means that the metallicity can now be used as the age proxy. Studying the behaviour of azimuthal velocity of the Milky Way stars as a function of [Fe/H], Belokurov and Kravtsov (2022) identify several key epochs in the life of the Galaxy, marked by noticeable changes in the stellar kinematics. For example, focusing on stars with retrograde motion, they see an edge to the Splash population at [Fe/H]$\approx-0.35$ in agreement with earlier studies (e.g. Bonaca et al., 2017; Haywood et al., 2018; Helmi et al., 2018; Di Matteo et al., 2019; Gallart et al., 2019; Belokurov et al., 2020). Across $-1<$[Fe/H]$<-0.35$, Splash co-exists with the high-$\alpha$ (“thick”) disc which, while heated by the interaction with the GS/E, appears largely intact, i.e. retains a relatively high amplitude of rotation. However, going further back in time to [Fe/H]$<-1$, the median azimuthal velocity drops precipitously low and the spread of the azimuthal velocity distribution reaches its largest value. In other words, starting instead from high redshift, the early Milky Way began without coherent rotation and then went through a rapid “spin-up” phase at $-1.3<$[Fe/H]$<-1$ from which point onward it has been dominated by a fast-rotating disc (see Fig. 18).
@@
To place their findings in context and reveal a plausible close-up view of the young Milky Way as predicted by current state-of-the-art models, Belokurov and Kravtsov (2022) compare the APOGEE+Gaia observations to numerical simulations of galaxy formation, namely FIRE (Wetzel et al., 2023) and Auriga (Grand et al., 2017). In both sets of simulations, Aurora (the state the Milky Way starts in) is characterised by a messy, turbulent spatial stellar distribution with a low overall angular momentum. Interestingly, some small amount of net prograde rotation is present at early epochs in most simulated Milky Way-like galaxies. Therefore, the in-situ population born before the disc appears to have angular momentum similar in amplitude and direction to the accreted halo (see Deason et al., 2017b). At the time of formation, Aurora stars can be seen close to the centre of the proto-Galaxy but lack order in both spatial and kinematic behaviour. As time passes and the Milky Way grows, Aurora settles and phase-mixes into a slightly flattened, quasi-spheroidal distribution with a very steep radial density profile. The Galactic spin-up phase is ubiquitous across the two simulation suites but the disc emergence typically happens at higher metallicities (and later times) compared to the observations. This view of the early Milky Way remains unchanged when other numerical simulation suites are considered, such as Illustris (Semenov et al., 2023) or ARTEMIS (Dillamore et al., 2023).
@@
Rix et al. (2022) use overall metallicity [M/H] inferred from the Gaia DR3 low-resolution XP spectro-photometry to provide a panoramic view of the low-metallicity Milky Way. They detect clear differences in the orbital and therefore spatial properties of their stars as a function of [M/H]. Above $-1.3<$[M/H]$<-1$, the Milky Way is dominated by a rapidly rotating disc, although accreted (GS/E) stars are also present. At [M/H]$<-1.3$ rotation quickly disappears (but does not go to zero) and the size of the orbital apo-centre shrinks: the bulk of the metal-poor population stays within the inner 2-3 kpc — these spatial and kinematic properties are consistent with the early studies of the inner Galaxy (Arentsen et al., 2020b, a). Rix et al. (2022) also detect a change in the slope of the metallicity distribution around [M/H]$\approx-1$ which they surmise must be connected with the onset of the disc formation 12 Gyr ago in agreement with age determinations of Xiang and Rix (2022). Rix et al. (2022) conclude that the Milky Way’s “poor old heart” that they see in the Gaia XP data is most likely the pre-disc proto-Galaxy and is the extension into the central Galactic regions of the same population which had been seen by Belokurov and Kravtsov (2022) and Conroy et al. (2022) in the Solar neighborhood.
@@ 4.3.2 Kraken, Koala, and Heracles
Figure 20: Left panel: The merger tree of the m12b simulation in Latte. The ‘proto-Milky Way’ is defined as the main branch halo (red) plus all the building blocks that merge (other colours) before $t_{\rm MR_{3:1}}$ (the time at which the main branch reaches a stellar mass ratio of 3:1 with the second most massive luminous subhalo). Right panel: Formation distance of star particles as a function of lookback time. At early times (i.e. before $t_{\rm MR_{3:1}}$) the main branch and its building blocks can have very similar chemical properties. [Reproduced from Horta et al. (2024).]
The total mass of the Milky Way’s stellar halo ($M_{\rm star}\approx 1-2\times 10^{9}M_{\odot}$, Deason et al., 2019; Mackereth and Bovy, 2020) puts a limit on the number of significant accretion events the Galaxy could have experienced. These mergers ought to be massive enough to sink deep in the Galactic potential and thus populate the Solar neighborhood with their tidal debris. Equally, their progenitor galaxies are likely amongst the main contributors to the accreted Milky Way GCs. The latter idea is explored in Kruijssen et al. (2019b), who compare the age-metallicity properties of the simulated and the observed Galactic GCs to infer the total number of large mass ratio accretion events in the Milky Way. Working with a suite of 25 zoom-in simulations, Kruijssen et al. (2019b) conclude that there have been at least two — but more likely three — main mergers in the Galaxy. The one with the lowest mass of these three is linked to the Sgr dwarf. The other two are more mysterious, in particular, the ancient interaction with the so-called ‘Kraken’ galaxy. As noticed quickly by Massari et al. (2019), the Kraken GCs identified by Kruijssen et al. (2019b) are not dynamically coherent and are likely from different progenitor systems. In divvying up the Galaxy’s GCs, Massari et al. (2019) find a group of 24 unassigned clusters at low total energy. These unclaimed low-energy GCs are classified by Forbes (2020) as a single progenitor named Koala, while Kruijssen et al. (2020) advocate that these GCs are part of Kraken. An alternative hypothesis is proposed by Belokurov and Kravtsov (2023a) and Belokurov and Kravtsov (2023b) who show that the bulk of the low-energy group originally highlighted by Massari et al. (2019) can instead be of in-situ origin and thus linked to the pre-disc Galactic population Aurora.
@@
Could instead a Kraken-like merger take place earlier in the history of the Galaxy, before its disc comes into existence? Horta et al. (2021) find chemo-dynamical evidence for such a primeval event in the central portion of the Milky Way using data from APOGEE and Gaia. Selecting stars with low [Al/Fe], indicative of the accreted population, they identify a gap in the total energy distribution, which separates the GS/E debris with high total energies from the more tightly bound, low-energy structure they dub ‘IGS’. They also give a name to the progenitor of the IGS — Heracles. Horta et al. (2021) demonstrate that the properties of the IGS are clearly distinct from those of the GS/E both in the chemical abundance space and in orbital space (although, as later demonstrated by Lane et al. 2022, the energy gap used for the IGS identification may instead be the imprint of the APOGEE selection function). However, differences between the IGS and the in-situ stars are somewhat less clear. Horta et al. (2021) show that at low metallicity and early epochs, the Milky Way in-situ chemical track passes through the region of low [Al/Fe], thus overlapping with the IGS selection. They conclude that the in-situ “bulge” population can have chemical properties identical to those of the IGS.
It is perhaps not impossible for the main progenitor of the Milky Way and a large, early-accreted dwarf galaxy to have similar chemistry if their masses were comparable. The dwarf’s star formation history would be truncated abruptly as it merged with the main progenitor of the Galaxy, but in the beginning (at lower metallicity), its chemical track and that of the Milky Way may look indistinguishable. This scenario is explored and quantified in Horta et al. (2024) where the FIRE simulations are used to compare the Milky Way’s main progenitor and its building blocks (see Fig. 20). The authors propose the following working definition of a “building block”: it is a dwarf galaxy that merges with the Milky Way during the period when the main progenitor is not yet clearly dominant within the simulation volume (typically around redshift $z\approx 3$). In the 13 galaxies within the FIRE suite, the main progenitor is easily identifiable: it contributes the most mass to the forming Milky Way, from $\sim 70\%$ to $\sim 95\%$. Horta et al. (2024) find that in 5 out of 13 galaxies considered, a building block exists that contributes between 15% and 30% of the total mass of the proto-Milky Way. While these building blocks have 2 to 5 times lower mass than the main progenitor, their chemical fingerprints are almost identical. Maybe, Heracles — the parent galaxy of the IGS — was such a building block and thus its chemistry and that of Aurora are indiscernible?
@@ 4.4 Implications of the GS/E and the early disc formation
An atypical assembly history? As we discussed in Section 1, the accretion of a small number of massive progenitors is not atypical for Milky Way-mass galaxies. However, what is perhaps more unusual in the case of the Milky Way is the lack of such events since the GS/E merger ($\sim 10$ Gyr ago). Evans et al. (2020) explored this explicitly and used the large-volume EAGLE simulations to quantify how ‘unusual’ the Milky Way accretion history is (at least with regard to massive progenitors). The authors find that only 5% of the Milky Way-mass haloes ($M_{200c}=0.7-2\times 10^{12}M_{\odot}$) in EAGLE have accreted a GS/E-like progenitor around 8-11 Gyr ago, but have had no subsequent massive merger events. This fraction lowers to $<1$ of haloes if the late-time accretion of the (surviving) LMC is also used as a constraint on the assembly history. An interesting consequence of the lack of other major mergers since the GS/E is that the haloes satisfying the Milky Way assembly history constraints in EAGLE are biased toward lower mass haloes within the ‘Milky Way-mass’ range, i.e. the haloes are typically $\leq 1\times 10^{12}M_{\odot}$. This bias is due to a long period of quiescence during which the haloes did not accrete as much mass as a ‘typical’ Milky Way galaxy.

### v2

@@ 4.2 Splash
The uncertainty in the precise timing of the events in the distant past of the Galaxy reflects the difficulty of obtaining reliable ages for large numbers of old stars, even in the Gaia era. Four main, distinct methods to infer stellar ages have been relied upon for archaeological studies, and each have been modernized and significantly improved thanks to the availability of the Gaia data. These four involve i) colour-magnitude diagram fitting, ii) astero-seismological mass estimates, iii) [C/N]-based mass estimates for red giants, and iv) isochronal fitting of the main sequence sub-giant and turn-off regions.
Gallart et al. (2019) demonstrate the power of the CMD fitting of large and homogeneous Gaia datasets to recover star-formation histories of the GS/E and the Splash in-situ stars. Montalbán et al. (2021) determine astero-seismological ages for a handful of stars belonging to the GS/E and the Splash using the data from the Kepler space observatory. Mackereth et al. (2019a) infer spectroscopic ages (relying on C and N abundances) for red giant stars in the APOGEE dataset using a neural network trained on astero-seismological age measurements and employ them to study the evolution of the disc heating with age. Sanders and Das (2018) measure isochronal ages for a combined catalogue of stars with spectroscopy from APOGEE, Gaia-ESO, GALAH, LAMOST, RAVE, and SEGUE.
Irrespective of the chronological method used, there appears to be a good deal of consensus as to the synchronicity of the SFHs of the GS/E and Splash. Both the accreted and the in-situ stars are mostly old, with relatively few stars younger than 8 Gyr. However, the exact shapes of the age distributions of the nearby halo stars do vary quite noticeably between the methods applied and the samples used (see Gallart et al., 2019; Bonaca et al., 2020; Belokurov et al., 2020; Xiang and Rix, 2022), and some works do find a tail of younger ages in the GS/E debris (e.g. Feuillet et al., 2021; Grunblatt et al., 2021; Horta et al., 2024b). It is currently not clear if the ages and/or association of the intermediate stars in the GS/E are robust, but, clearly accurate ages are a vital ingredient for future studies of early-time Galactic archaeology.
@@ 4.3 Pre-disc Milky Way
Luckily, additional chemical tags can be used to attempt to separate the accreted and in-situ born stars at [Fe/H]$<-1$. These ideas are first presented in Hawkins et al. (2015) who suggest that the accreted and in-situ halo stars attain distinct levels of [Al/Fe] abundance ratio. The stellar nucleosynthetic yield of aluminium (and sodium) is low at low metallicity but increases dramatically as more of C, N, O is pumped into the interstellar medium (see Kobayashi et al., 2006) which causes a delay in Al enrichment compared to conventional $\alpha$-elements. Compared to more massive systems, such as the progenitor of the Milky Way, dwarf galaxies are less efficient at retaining and re-using gas during star formation and are therefore slower to reach higher levels of metallicity when Al production becomes efficient. As a result, in a dwarf, by the time Al starts to be produced in earnest, type Ia SNe begin to dominate Fe production, thus inhibiting the growth of the [Al/Fe] ratio. In a more rapidly evolving Milky Way progenitor, Al production kicks in before the Type Ia SNe contribution rises and thus the young Galaxy can enjoy a period of over-abundance of Al relative to Fe. The systematic difference between [Al/Fe] levels in dwarf galaxies and the Milky Way proper is now well established (see e.g. Hasselquist et al., 2021) and is routinely taken advantage of to identify accreted stars in the nearby halo (e.g. Hawkins et al., 2015; Das et al., 2020; Horta et al., 2021).
@@ 4.3.1 Aurora
Instead of focusing on the accreted debris, Belokurov and Kravtsov (2022) combine [Al/Fe] measurements published as part of the APOGEE DR17 and the Gaia EDR3 astrometry to create a pure Milky Way in-situ sample, including stars at low metallicities (see top panel of Fig. 18). They show that while there may be some amount of false negatives, i.e. genuine in-situ stars identified as accreted, the in-situ selection is largely uncontaminated. Having access to a pure set of stars born only in the Milky Way means that the metallicity can now be used as the age proxy. Studying the behaviour of azimuthal velocity of the Milky Way stars as a function of [Fe/H], Belokurov and Kravtsov (2022) identify several key epochs in the life of the Galaxy, marked by noticeable changes in the stellar kinematics. For example, focusing on stars with retrograde motion, they see an edge to the Splash population at [Fe/H]$\approx-0.35$ in agreement with earlier studies (e.g. Bonaca et al., 2017; Haywood et al., 2018; Helmi et al., 2018; Di Matteo et al., 2019; Gallart et al., 2019; Belokurov et al., 2020). Across $-1<$[Fe/H]$<-0.35$, Splash co-exists with the high-$\alpha$ (“thick”) disc which, while heated by the interaction with the GS/E, appears largely intact, i.e. it retains a relatively high amplitude of rotation. However, going further back in time to [Fe/H]$<-1$, the median azimuthal velocity drops precipitously low and the spread of the azimuthal velocity distribution reaches its largest value. In other words, starting instead from high redshift, the early Milky Way began without coherent rotation and then went through a rapid “spin-up” phase at $-1.3<$[Fe/H]$<-1$ from which point onward it has been dominated by a fast-rotating disc (see Fig. 18).
@@
To place their findings in context and reveal a plausible close-up view of the young Milky Way as predicted by current state-of-the-art models, Belokurov and Kravtsov (2022) compare the APOGEE+Gaia observations to numerical simulations of galaxy formation, namely FIRE (Wetzel et al., 2023) and Auriga (Grand et al., 2017). In both sets of simulations, Aurora (the state the Milky Way starts in) is characterized by a messy, turbulent spatial stellar distribution with a low overall angular momentum. Interestingly, some small amount of net prograde rotation is present at early epochs in most simulated Milky Way-like galaxies. Therefore, the in-situ population born before the disc appears to have angular momentum similar in amplitude and direction to the accreted halo (see Deason et al., 2017b). At the time of formation, Aurora stars can be seen close to the centre of the proto-Galaxy but lack order in both spatial and kinematic behaviour. As time passes and the Milky Way grows, Aurora settles and phase-mixes into a slightly flattened, quasi-spheroidal distribution with a very steep radial density profile. The Galactic spin-up phase is ubiquitous across the two simulation suites but the disc emergence typically happens at higher metallicities (and later times) compared to the observations. This view of the early Milky Way remains unchanged when other numerical simulation suites are considered, such as Illustris (Semenov et al., 2024) or ARTEMIS (Dillamore et al., 2023b). As a result of the Aurora discourse, recent works have paid more specific attention to the proto-disc galaxy populations in cosmological simulations (see e.g. Horta et al., 2024a; McCluskey et al., 2024).
@@
Rix et al. (2022) use overall metallicity [M/H] inferred from the Gaia DR3 low-resolution XP spectro-photometry to provide a panoramic view of the low-metallicity Milky Way. Importantly, this work is not limited to the solar neighourhood, which is the vicinity where Aurora was originally discovered. Rix et al. (2022) detect clear differences in the orbital and therefore spatial properties of their stars as a function of [M/H]. Above $-1.3<$[M/H]$<-1$, the Milky Way is dominated by a rapidly rotating disc, although accreted (GS/E) stars are also present. At [M/H]$<-1.3$ the rotation quickly disappears and eventually goes to zero below [Fe/H]$=-2$. Rix et al. (2022) also detect a change in the slope of the metallicity distribution around [M/H]$\approx-1$ which they surmise must be connected with the onset of the disc formation 12 Gyr ago in agreement with age determinations of Xiang and Rix (2022). Rix et al. (2022) conclude that the Milky Way’s “poor old heart” that they see in the Gaia XP data is most likely the pre-disc proto-Galaxy and is the extension into the central Galactic regions of the same population which had been seen by Belokurov and Kravtsov (2022) and Conroy et al. (2022) in the Solar neighourhood.
A clearer picture of the orbital structure in the inner metal-poor halo emerges in the analysis of the Pristine Inner Galaxy Survey (PIGS) data (Arentsen et al., 2020b, a). As Ardern-Arentsen et al. (2024) demonstrate, below [Fe/H]$\approx-1$ the typical size of the orbital apo-centre is small, $<5$ kpc. However, at lower metallicities, i.e. [Fe/H]$<-2$ it starts to grow: only $\approx 60\%$ of stars remain confined within 5 kpc. Thus, while Ardern-Arentsen et al. (2024) agree that the Milky Way hosts a significant, centrally concentrated, metal-poor component with a modest prograde rotation, likely connected to the pre-disc proto Galaxy (i.e. Aurora), they also see clear evidence of a more extended halo component in the very metal-poor regime, i.e. at [Fe/H]$<-2$, which may be of accreted nature. The above PIGS study concurs with the earlier work of Lucey et al. (2021) who saw a similar pattern in the change of the apo-centre with metallicity in their smaller COMBS survey sample. Larger samples of (very) metal-poor stars from upcoming spectroscopic surveys (e.g. 4MOST) are on the horizon, and thus there is great scope to shed further light on the earliest stages of the Galaxy’s evolution in the coming years.
@@ 4.3.2 Kraken, Koala, and Heracles
Figure 20: Left panel: The merger tree of the m12b simulation in Latte. The ‘proto-Milky Way’ is defined as the main branch halo (red) plus all the building blocks that merge (other colours) before $t_{\rm MR_{3:1}}$ (the time at which the main branch reaches a stellar mass ratio of 3:1 with the second most massive luminous subhalo). Right panel: Formation distance of star particles as a function of lookback time. At early times (i.e. before $t_{\rm MR_{3:1}}$) the main branch and its building blocks can have very similar chemical properties. [Reproduced from Horta et al. (2024a).]
The total mass of the Milky Way’s stellar halo ($M_{\rm star}\approx 1-2\times 10^{9}M_{\odot}$, Deason et al., 2019; Mackereth and Bovy, 2020) puts a limit on the number of significant accretion events the Galaxy could have experienced. These mergers ought to be massive enough to sink deep in the Galactic potential and thus populate the Solar neighourhood with their tidal debris. Equally, their progenitor galaxies are likely amongst the main contributors to the accreted Milky Way GCs. The latter idea is explored in Kruijssen et al. (2019b), who compare the age-metallicity properties of the simulated and the observed Galactic GCs to infer the total number of large mass ratio accretion events in the Milky Way. Working with a suite of 25 zoom-in simulations, Kruijssen et al. (2019b) conclude that there have been at least two — but more likely three — main mergers in the Galaxy. The one with the lowest mass of these three is linked to the Sgr dwarf. The other two are more mysterious, in particular, the ancient interaction with the so-called ‘Kraken’ galaxy. As noticed quickly by Massari et al. (2019), the Kraken GCs identified by Kruijssen et al. (2019b) are not dynamically coherent and could therefore originate from different progenitor systems. In dividing up the Galaxy’s GCs, Massari et al. (2019) find a group of 24 unassigned clusters at low total energy. These unclaimed low-energy GCs are classified by Forbes (2020) as a single progenitor named Koala, while Kruijssen et al. (2020) advocate that these GCs are part of Kraken. An alternative hypothesis is proposed by Belokurov and Kravtsov (2024) and Belokurov and Kravtsov (2023) who show that the bulk of the low-energy group originally highlighted by Massari et al. (2019) can instead be of in-situ origin and thus linked to the pre-disc Galactic population Aurora.
@@
Could instead a Kraken-like merger take place earlier in the history of the Galaxy, before its disc comes into existence? Horta et al. (2021) find chemo-dynamical evidence for such a primeval event in the central portion of the Milky Way using data from APOGEE and Gaia. Importantly, this work (unlike others which are limited to the solar neighourhood) investigates the nature of halo stars within 5 kpc from the Galactic centre, and thus, presumbably, in the vicinity where the majority of the mass from proto-Galaxy populations reside. Selecting stars with low [Al/Fe], indicative of the accreted population, they identify a gap in the total energy distribution, which separates the GS/E debris with high total energies from the more tightly bound, low-energy structure they dub ‘IGS’. They also give a name to the progenitor of the IGS — Heracles. Horta et al. (2021) demonstrate that the properties of the IGS are clearly distinct from those of the GS/E both in the chemical abundance space and in orbital space (although, as later demonstrated by Lane et al. 2022, the energy gap used for the IGS identification may instead be the imprint of the APOGEE selection function). However, differences between the IGS and the in-situ stars are somewhat less clear. Horta et al. (2021) show that at low metallicity and early epochs, the Milky Way in-situ chemical track passes through the region of low [Al/Fe], thus overlapping with the IGS selection. They conclude that the in-situ “bulge” population can have chemical properties identical to those of the IGS.
It is perhaps not impossible for the main progenitor of the Milky Way and a large, early-accreted dwarf galaxy to have similar chemistry if their masses were comparable. The dwarf’s star formation history would be truncated abruptly as it merged with the main progenitor of the Galaxy, but in the beginning (at lower metallicity), its chemical track and that of the Milky Way may look indistinguishable. This scenario is explored and quantified in Horta et al. (2024a) where the FIRE simulations are used to compare the Milky Way’s main progenitor and its building blocks (see Fig. 20). The authors propose the following working definition of a “building block”: it is a dwarf galaxy that merges with the Milky Way during the period when the main progenitor is not yet clearly dominant within the simulation volume (typically around redshift $z\approx 3$). In the 13 galaxies within the FIRE suite, the main progenitor is easily identifiable: it contributes the most mass to the forming Milky Way, from $\sim 70\%$ to $\sim 95\%$. Horta et al. (2024a) find that in 5 out of 13 galaxies considered, a building block exists that contributes between 15% and 30% of the total mass of the proto-Milky Way. While these building blocks have 2 to 5 times lower mass than the main progenitor, their chemical fingerprints are almost identical. Maybe, Heracles — the parent galaxy of the IGS — was such a building block and thus its chemistry and that of Aurora are indiscernible?
@@ 4.4 Implications of the GS/E and the early disc formation
An atypical assembly history? As we discussed in Section 1, the accretion of a small number of massive progenitors is not atypical for Milky Way-mass galaxies. However, what is perhaps more unusual in the case of the Milky Way is the lack of such events since the GS/E merger ($\sim 10$ Gyr ago). Evans et al. (2020) explored this explicitly and used the large-volume EAGLE simulations to quantify how ‘unusual’ the Milky Way accretion history is (at least with regard to massive progenitors). The authors find that only 5% of the Milky Way-mass haloes ($M_{200c}=0.7-2\times 10^{12}M_{\odot}$) in EAGLE have accreted a GS/E-like progenitor around 8-11 Gyr ago, but have had no subsequent massive merger events. This fraction lowers to $<1$ of haloes if the late-time accretion of the (surviving) LMC is also used as a constraint on the assembly history. An interesting consequence of the lack of other major mergers since the GS/E is that the haloes satisfying the Milky Way assembly history constraints in EAGLE are biased towards lower mass haloes within the ‘Milky Way-mass’ range, i.e. the haloes are typically $\leq 1\times 10^{12}M_{\odot}$. This bias is due to a long period of quiescence during which the haloes did not accrete as much mass as a ‘typical’ Milky Way galaxy.